OpenStudy (anonymous):

can someone please solve -(1/6)+(x/2)(16+x^2)^(-1/2)=0 for x?

OpenStudy (anonymous):

Ok, I got it this time, sorry for wrong solution posted initially. This equation can be rewritten as \[\frac{ 1 }{ 6} = \frac{ x }{ 2\sqrt{16 + x^2} }\] Then Multiply both sides by denominators to eliminate fractions. This results in \[2\sqrt{16 + x^2} = 6x\] Divide both sides by two to isolate the radical, then square both sides to eliminate the radical. This results in \[16 + x^2 = 9x^2 \] Then subtract both sides by x^2 and then solve for x (I will leave these rudimentary steps for you to work on!). The final answer should be \[x =\pm \sqrt{2}\]